OpenStudy (lolaus):

Solve the equation. 2^3square root 5m+5 +7=5

OpenStudy (wckeller):

Can you use the equation button below to show your problem? It's kind of difficult to understand where the square root stops.

OpenStudy (lolaus):

\[2^{3}\sqrt{5m+4}+7=5\]

OpenStudy (wckeller):

is it 5m+4 or 5m+5 because in the first representation of the equation you said it was a +5

OpenStudy (wckeller):

Oh! By the way, welcome to Open Study!

OpenStudy (lolaus):

my bad the right question is 5m+4

OpenStudy (wckeller):

Okay. So to solve for M we need to get it all alone on one side of the equation. First we can start by subtracting a 7 from both sides.

OpenStudy (lolaus):

thank you appreciated.

OpenStudy (wckeller):

\[2^3 \sqrt{5m+4}=5-7\]

OpenStudy (wckeller):

\[2^3 \sqrt{5m+4}=-2\]

OpenStudy (wckeller):

\[2^3 = 8\] \[8\sqrt{5m+4}=-2\]

OpenStudy (wckeller):

Divide both sides by 8

OpenStudy (wckeller):

\[\sqrt{5m+4}=\frac{ -2 }{ 8 }\]

OpenStudy (wckeller):

Now in order to get rid of the square root, we need to square both sides.

OpenStudy (wckeller):

\[\sqrt{5m+4}^2=(\frac{ -2 }{ 8 })^2\] \[5m+4=\frac{ 4 }{ 64 }\] \[\frac{ 4 }{ 64 } = \frac{ 1 }{ 16 }\] \[5m+4=\frac{ 1 }{ 16 }\]

OpenStudy (wckeller):

Now we need to subtract 4 from both sides.

OpenStudy (wckeller):

\[5m=\frac{ 1 }{ 16 }-4\] Find a common denominator for 4 and 1/16 \[4=\frac{ 64 }{ 16 }\] \[5m=\frac{ 1 }{ 16 }-\frac{ 64 }{ 16 }\]

OpenStudy (wckeller):

\[1-64=-63\] \[5m=\frac{ -63 }{ 16 }\] Now we need to divide both sides by 5. \[m=\frac{ \frac{ -63 }{ 16 } }{ 5 }\] Which is the same as \[m=\frac{ -63 }{ 16*5 }\] \[m=\frac{ -63 }{ 80 }\] Final answer: \[m=\frac{ -63 }{ 80 }\]

OpenStudy (wckeller):

I may have messed up somewhere, let me plug back in and check my work.

OpenStudy (lolaus):

Thank you Wckeller I really appreciate your kind help.

OpenStudy (wckeller):

Okay, I am SO sorry... I have royally messed up somewhere.

OpenStudy (wckeller):

The answer I'm getting when I plug back in is 9.

OpenStudy (wckeller):

But its supposed to be 5... let me find where I messed up.

OpenStudy (wckeller):

Oh boy, ive really messed up somewhere

OpenStudy (wckeller):

I found it.. So there is NO answer to this problem. A square root cannot equal a negative number.

OpenStudy (wckeller):

\[\sqrt{5m+4}=\frac{ -2 }{ 8 }\] This is where the problem occurs.

OpenStudy (wckeller):

So the answer to this problem is: No Solution